3Fn − Fn−2 = Fn+2, for n ≥ 3.


1
Expert's answer
2021-12-08T12:59:14-0500

The sequence of Fibonacci numbers, "F_0, F_1, F_2,F_3, . . .," are defined by the following equations: 


"F_0=0,""F_1=1,""F_n=F_{n-1}+F_{n-2}, n\geq 2"

Then for "n\geq 3"

"F_{n+2}=F_{n+1}+F_n=F_n+F_{n-1}+F_n"

"=2F_n+F_n-F_{n-2}=3F_n-F_{n-2}"

Therefore


"3F_n-F_{n-2}=F_{n+2}, n\geq 3"